OpenStudy (anonymous):

-Multiply. (–4) ∙ (–8) A. –32 B. –12 C. 12 D. 32

OpenStudy (anonymous):

@sweetburger @sweetburger

OpenStudy (midhun.madhu1987):

-4 * -8 = - * - = + 4 * 8 = 32 So what is the answer?

OpenStudy (anonymous):

32... when you mutiply 2 negatives together it always is positive, always check your signs

OpenStudy (anonymous):

It's like saying 4 * 8. When you multiply two negatives, it will always be positive, so the answer is 32. Hope that helps!

OpenStudy (anonymous):

What is 5 • (–13)?

OpenStudy (anonymous):

What is 5 • (–13)?

OpenStudy (anonymous):

-65

OpenStudy (anonymous):

5 times (-13)

OpenStudy (anonymous):

yeah.. -65

OpenStudy (anonymous):

Multiply. -1.1 times (-8) =

OpenStudy (anonymous):

@Cman456

OpenStudy (anonymous):

Multiply. (–2.1) ∙ (–1.4) A. –3.5 B. –2.94 C. 2.94 D. 3.5

OpenStudy (anonymous):

multiply (-2.1) times (-1.4)=

OpenStudy (anonymous):

I hope i get a metal for this..

OpenStudy (anonymous):

-1.1x-8 = 8.8...... -2.1x-1.4=2.94...